I am trying to find the slope of a line that goes through the origin and the point(10,5) without graphing. I don't know what the origin is so I am very lost with this problem. I thought I could solve it using y-y1=m(x-x1) but it doesn't work. Please advise.
Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39617)

Do not try to make this complicated. Start from the point (0,0), and find the vertical change and horizontal change, and form the ratio of them.

Answer by stanbon(75887)

I am trying to find the slope of a line that goes through the origin and the point(10,5) without graphing. I don't know what the origin is so I am very lost with this problem. I thought I could solve it using y-y1=m(x-x1) but it doesn't work. Please advise.
----
Note: The origin is the point (0,0)
-----
slope = m = (y-y1)/(x-x1) = (5-0)/(10-0) = 5/10 = 1/2
-----------
Cheers,
Stan H.
-----------